OpenStudy (anonymous):

find 18 "sigma notation" i=7 11i+3

OpenStudy (anonymous):

18 is the upper limit i=7 is the lover limit and 11i+3 is next to the sigma notation

OpenStudy (amistre64):

i changes from 7 to 18 11 (7) + 3 +11 (8) + 3 +11(9) + 3 +11(10) + 3 +11(11) + 3 ....

OpenStudy (amistre64):

sum (from i=7 to 18) of f(i) is a good way to describe it

OpenStudy (anonymous):

so I do that till i get to 18?

OpenStudy (anonymous):

what do I do after that?

OpenStudy (amistre64):

that is the brute math way; yes

OpenStudy (amistre64):

there are more elegant methods im sure

OpenStudy (amistre64):

\[\sum_{i=7}^{18}11i+\sum_{i=7-6}^{18-6}3\]

OpenStudy (anonymous):

so wait Im trying to find the sum of all those numbers? So do I add them?

OpenStudy (amistre64):

yes, sum means to add up all those values

OpenStudy (anonymous):

\[n/2(a+l)\]

OpenStudy (amistre64):

\[11\sum_{i=7}^{18}i+3\sum_{i=1}^{12}\] \[11\frac{12(7+18)}{2}+3(12)\]\[66(25)+36\] \[1650+36=1686\] if i did it right :)

OpenStudy (anonymous):

oh my gosh thankyou soo much! Thats an answer choice, and I think i get it a little better! You're the best!

OpenStudy (amistre64):

just because its an answer choice, doesnt mean its right .... but yes, its right lol

OpenStudy (anonymous):

interesting...

OpenStudy (anonymous):

@amistre, now you have me thinking about how to prove:\[a_{1}+a_{2}...a_{n}=n \frac{(a_{n}+a_{1})}{2}\]in general.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Ok, after pondering this one, it isn't true in general. It is true when the terms of the series are related linearly.

OpenStudy (amistre64):

:) its true of an arithmetic progression ... if im remembering correctly

OpenStudy (amistre64):

or as you put it, a linear relationship ... yes

OpenStudy (anonymous):

Yeah, i'm very rusty with sequences and series...still plugging away at this idea :)

OpenStudy (anonymous):

what is a linear relationship?

OpenStudy (anonymous):

A linear relationship has the form:\[y=mx+b\] where m and b are constants.